OpenStudy (anonymous):

Find the antilogarithm of the logarithm: LOG a= 0.7380 LOG g= 0.6990

OpenStudy (anonymous):

Does your logarithm have an assumed base of 10 here? If so, by the definition of the logarithm, \[\log a = 0.7380 \iff 10^0.7380 = a\]and\[\log g = 0.6990 \iff 10^0.6990 = g\]

OpenStudy (anonymous):

g would be approximately equal to 5 a would be approximately equal to 5.47